The table below displays an economy that produces only two goods: rubber ducks and laptop chargers for the years 2015 and 2016. Using the GDP-deflator, the inflation rate between 2015 and 2016 is approximately equal to

2015

price./quantity

2016

price/quantity

rubber ducks $2 / 3,000 $3 / 2,000
laptops $10 / 4,000 $11.50 / 3,000

13.58%

1.358%

19.12%

1.912%

100%

1%
Answer #1

Ans: 19.12%

Explanation:

Nominal GDP in 2015 = ($2 * 3000) + ($10 * 4000) = $46,000

Real GDP in 2015 = ($2 * 3000) + ($10 * 4000) = $46,000

Since 2015 is base year, Nominal GDP and Real GDP are equal

GDP Deflator = (Nominal GDP / Real GDP) * 100

                      = ($46,000 / $46,000) * 100

                      = 100

Nominal GDP in 2016 = ($3 * 2000) + ($11.50 * 3000) = $40,500

Real GDP in 2016 = ($2 * 2000) + ($10 * 3000) = $34,000           

GDP Deflator = (Nominal GDP / Real GDP) * 100

                      = ($40,500 / $34,000) * 100

                      = 119.12

The inflation rate between 2015 and 2016 = [(119.12 - 100) / 100] * 100 = 19.12%